Find a 2×2 singular matrix A that maps
(1, 1)^T into (1, 3)^T
1
Expert's answer
2020-12-21T17:32:51-0500

Let "A=\left[\begin{array}{cc} a & b \\ c & d\end{array}\right]" is a singular matrix. Then "\det A=ad-bc=0". We have that


"\left[\begin{array}{cc} a & b \\ c & d\end{array}\right] \cdot \left[\begin{array}{c} 1 \\ 1\end{array}\right]=\left[\begin{array}{c} 1 \\ 3\end{array}\right]" and thus, "a+b=1" and "c+d=3". It follows that "a=1-b" and


"d=3-c", and therefore, "0=ad-bc=(1-b)(3-c)-bc=3-3b-c+bc-bc=3-3b-c", and thus "c=3-3b" .


Let "b=0". Then "c=3,\ a=1-b=1,\ d=3-c=3-3=0."


We conclude that the matrix "A=\left[\begin{array}{cc} 1 & 0 \\ 3 & 0\end{array}\right]" is  singular that maps "(1, 1)^T" into "(1, 3)^T".